| Entry | Date | Title | Site | Author | #Graphics | 
|---|---|---|---|---|---|
| 189 | Tue 04-Mar-2003 | Gravimetric soil samples | all stns | oncley | |
| 164 | Mon 24-Feb-2003 | Deep soil moisture at Lake | lake | delany | |
| 159 | Sat 22-Feb-2003 | 10V regulators added | all stns | oncley | |
| 152 | Thu 20-Feb-2003 | Deep soil moisture at Sage | sage | delany | |
| 128 | Tue 28-Jan-2003 | Sage / Lake soil cable problem checks | lake | militzer | |
| 123 | Tue 28-Jan-2003 | Grass TP01 soil properties fixed | tower | militzer | |
| 118 | Wed 22-Jan-2003 | Soil sensor configuration | all stns | delany | 1 | 
| 116 | Wed 22-Jan-2003 | TP01 wiring | all stns | semmer | 1 | 
| 101 | Tue 07-Jan-2003 | testing of thermal properties probes | all stns | maclean | |
| 84 | Thu 19-Dec-2002 | Corrected therm prop. timing at lake | lake | maclean | |
| 80 | Wed 18-Dec-2002 | Soil gravimetric/bulk density | all stns | delany | |
| 70 | Tue 17-Dec-2002 | Gravimetric soil samples | all stns | delany | |
| 67 | Mon 16-Dec-2002 | Gravimetric soil samples | all stns | delany | |
| 61 | Mon 16-Dec-2002 | TP01 installed | lake | oncley | |
| 60 | Mon 16-Dec-2002 | TP01 installed | tower | oncley | |
| 54 | Thu 12-Dec-2002 | Changed Tsoil sampling at sage,lake | lake | maclean | |
| 51 | Wed 11-Dec-2002 | Modifed grass logger Tsoil sampling | tower | maclean | |
| 46 | Mon 09-Dec-2002 | soil profile array cable installed | tower | knudson | |
| 7 | Mon 14-Oct-2002 | FLOSS02 soil sensors | all stns | delany | |
| 4 | Fri 11-Oct-2002 | Hft Probes 1,2 upside down | lake | militzer | |
| 3 | Fri 11-Oct-2002 | Soil Echo Probes Rotated to Vertical | all stns | militzer | |
| 1 | Tue 08-Oct-2002 | Installation of soil sensors | all stns | delany | 
17:00 Mar 4 Sage
Near B.
                   Fresh(4 Mar) Dry(5 Mar)	%(wet/total)
A sample taken 16:00 which broke, but I recovered some:
   0-~5	   	   80.3		56.9		29.1
		   56.9		49.7		12.7
The real sample:
   0-3 cm  EI 	   28.0	    	18.4 		34.3
           EII     30.0		19.8		34.0
           EIII    28.7		19.8		31.0
   Mean:					33.1
   Total sample:   88.6         [59.3]
   Density:			[0.86]
   3-8 cm  FI      52.8		44.1		16.5
           FII     54.0		44.5		17.6
           FIII    48.8		38.4		21.3
   Mean:					18.5
   Total sample:  179.3		[146.1]
   Density:			[1.28]
12:45 Mar 4 Lake
Close to B!
Dug down through 28 cm of snow - broke up ice coating surface
                   Fresh(4 Mar) Dry(5 Mar)
(has some ice/snow on the top, but mostly solid frozen)
  0-3 cm  AI       38.0		28.5		25.0
          AII      35.9		27.3		24.0
          AIII     23.3		17.8		23.6
  Mean:						24.2
  Total sample:	   99.8		[75.6]
  Density			[1.10]
(peaty: root and stem)
  3-8 cm  BI       19.3		14.7		23.8
          BII      15.7		11.3		28.0
          BIII     16.0		11.2		30.0
  Mean:						27.3	
  Total sample:    87.8         [63.8]
  Density:	   		[0.56]
15:00 Mar 5 Grass site:
Close to B:
Dug down through 28 cm of snow - broke up ice coating surface
                   Fresh(5 Mar) Dry(6 Mar)
(has some ice/snow on the top, but mostly solid frozen)
  0-3 cm  CI       18.3		4.4		76.0
          CII      10.9		3.1		71.6
          CIII     12.9		2.9		77.5
  Mean:						[75.0]
  Total sample:	   42.9		[10.7]
  Density	   0.62		0.16
(peaty: root and stem)
  3-8 cm  DI       28.3		22.1		21.9
          DII      19.8		13.1		33.8
          DIII     29.4		24.0		18.4
  Mean:						[24.7]
  Total sample:    111.2	[83.7]
  Density:	   0.97		0.73
17:00 Mar 5 Lake site:
Close to C:
                   Fresh(5 Mar) Dry(
(has some ice/snow on the top, but mostly solid frozen)
  0-3 cm  EI       32.8		23.5		28.4	[had a grub inside!]
          EII      26.9		19.7		26.8
          EIII     28.4		20.6		27.5
  Mean:						[27.5]
  Total sample:	   92.5		[67.1]
  Density	   1.35		0.98
(peaty: root and stem)
  3-8 cm  FI       39.4		24.7		37.3
          FII      44.8		32.3		27.9
          FIII     39.8		22.7		43.0
  Mean:						[36.1]
  Total sample:    157.9	[100.9]
  Density:	   1.38		0.88
09:30 Mar 6 Grass site:
Close to A:
                   Fresh(6 Mar) Dry(7 Mar)
(has some ice/snow on the top, but mostly solid frozen)
  0-3 cm  AI       14.5		1.6		89.0
          AII      10.6		1.0		90.6
          AIII     11.1		1.1		90.1
						89.8
  Mean:			
  Total sample:	   37.0 	3.8		(but not complete volume)
  Density	   0.54		0.06
(peaty: root and stem)
  3-8 cm  BI       22.2		9.8		55.9
          BII      16.5		7.5		54.5
          BIII     17.9		4.6		74.3
  Mean:						61.3
  Total sample:    87.8		34.0
  Density:	   0.77		0.30
13:00 Mar 6 Sage site:
Close to A:
I took two cores that broke (along with the corer), so I only recovered about 
6cm deep cores twice.
                   Fresh(6 Mar) Dry(7 Mar)
  0-~6 cm EI       44.2		34.3		22.4
          EII      34.5		27.3		20.9
          EIII     42.4		34.0		19.8
  Mean:						21.1
  Total sample:	   138.7 	109.4		(very crude estimate)
  Density	   1.01		0.80
  0-~6 cm FI       42.8		34.5		19.4
          FII      37.7		31.4		16.7
          FIII     39.4		31.8		19.3
  Mean:						18.5
  Total sample:    168.1 	137.0		(very crude estimate)
  Density:	   1.22		1.00
[] values are calculated from other variables
Diameter of corer is 5.4cm
At Lake site at 15:00- 16:00. Level with the middle soil installations, 6 - 7m
north of the center of the radiometer stand.
Cleared the 15 cm snow for a 1m diameter circle. 
There was an ice crust a cm or so thick. This was broken up and discarded.
Then carefully excluding
loose snow dug a 20 cm wide, 35 cm deep pit using a digging bar. The loose dirt
was sccoped out by hand and thrown to the north, away from the rad stand.
Using trowel took samples of 0-5cm,5-10cm,10-15cm,15-20cm,20-25cm,and 25-30 cm 
Returned to base and weighed aliquots
Into oven 16:40 on 24 Feb''03
Out of oven for weighing 09:00 on 26 Feb''03
         Aliquots   Fresh       Dry      Different    Gravmoist
 0-5cm    A I       33.4        25.4      8.0        0.315   
          A II      36.1        27.5      8.6        0.313     30.4%
          AIII      30.6        23.8      6.8        0.286
 5-10cm   BI        28.4        23.7      4.7        0.198
          BII       26.5        22.2      4.3        0.194     20.1%
          BIII      27.5        22.7      4.8        0.211
 10-15cm  CI        26.4        23.0      3.4        0.148
          CII       29.1        25.4      3.7        0.146     14.9%
          CIII      25.8        22.4      3.4        0.152
 15-20cm  DI        24.0        20.7      3.3        0.159
          DII       23.0        19.7      3.3        0.168     16.4%
          DIII      22.0        18.9      3.1        0.164
 20-25cm  EI        28.3        23.8      4.5        0.189
          EII       26.2        22.0      4.2        0.191     18.9%
          EIII      24.1        20.3      3.8        0.187
 25-30cm  FI        30.6        25.1      5.5        0.219
          FII       24.7        20.3      4.4        0.217     21.8%
          FIII      24.1        19.8      4.3        0.217
There is a minimum soil moisture at ~ 15 cm. Note that below this depth the soil
changes from "sandy" to "clayey".
From 1300-1400 added regulators to the switched 12V signals from the CR10s at all 3 sites (grass, then sage, then lake). I'm not sure about some of the electrical contacts -- the daily plots show that I did make a change -- hopefully, it was for the good!
To investigate the soil moisture down to the 35 cm depth, a pit was dug at the
Sage site. The pit was 5m or so south of the Darkhorse. near the fence.
It was in a small clear patch with 20-30cm sage bushes some 20-30 cm away.
Several cm of snow was brushed away and with digging bars the ground was broken
to a depth, then the loose dirt collected.
Samples were collected for:
                            0 -  6 cm
                            6 - 10 cm
                           10 - 15 cm
                           15 - 20 cm
                           20 - 25 cm
                           30 - 35 cm
Triplicate aliquots were separated and introduced into the oven at 12:00
The dried aliquots were removed and weighed the next day at 09:45.
Note that the ordering of the aliquots C, B, A, D, E, F is correct
   Sample    Aliquot   Fresh      Dry     Difference      Gravmoist
                         g         g          g
  0 -  6 cm   C I      22.7       19.0       3.7        0.195
              C II     27.5       23.4       4.1        0.175   18.7%
              C III    25.6       21.5       4.1        0.191
  6 - 10 cm   B I      21.8       19.0       2.8        0.147
              B II     18.2       15.8       2.4        0.152   15.0%
              B III    24.3       21.1       3.2
 10 - 15 cm   A I      24.5       21.5       3.0        0.140
              A II     23.0       20.3       2.7        0.152   13.6%
              A III    20.9       18.4       2.5        0.152
 15 - 20 cm   D I      28.0       24.7       3.3        0.134
              D II     25.5       22.5       3.0        0.133   13.2%
              D III    26.1       23.1       3.0        0.130
 20 - 25 cm   E I      21.3       19.0       2.3        0.121
              E II     24.0       21.3       2.7        0.127   12.7%
              E III    24.7       22.0       2.7        0.123
 30 - 35 cm   F I      19.9       17.9       2.0        0.112
              F II     26.2       23.6       2.6        0.110   11.2%
              F III    25.7       23.1       2.6        0.113
7/28/03 we examined the soil problems noted in the 'todo' list. SAGE: Very little snow on ground, mostly bare with patches of snow. gsoil.c.10cm yes indeed the cable was spliced and it appears dicey to try to fix it where it enters the frozen ground. Tsoil.a.2.5cm appeared ok and not disturbed even though someone said it looks 'wierd' took pictures. LAKE: crusted snow cover all over. Perhaps 4-8" deep. Gsoil.c.10cm has a wiring break and we were supposed to resolder it. We played a bit trying to pull out the cable to find the 'tape job' but decided we were impacting the surface and jeopardizing the other sensors so stopped. Qsiol.c.10cm ditto. Logger connections are secure and look good.
7/28/03 ~1pm TP01 should now be up. We haven't seen a pulse on the Vheat yet but it's still set to 3-hour cycling and we didn't change that. Checked wiring after examining sage site which is working. CR10 wiring OK. I did note the C7 - Switched 12 goes to the access point 2 positions away instead of the access across the terminal blocks adjacent to the +12 input. Amp-9 connector from probe was bad. It could never have worked and apparently wasn't checked in Boulder before coming up here. Pin 2 (gnd) was crimped with the insulation preventing contact. Pin 1/4 were in wrong positions. Pin 8 with 150 ohm connector had and has a tenuous solder to the wire. We probably broke off when unscrewing the backshell. Wiring correct now. NOTE: the 150ohm connection needs to be redone so it's more secure.
To view the configuration of the soil sensors at Grass, Sage and Lake: Hit "Graphics Viewer", and then the little diamond.
The graphics file shows the wiring ofr the TP01 probe. The color code for the wires corrresponds to the probe wires and NOT the wires in the logger box.
At sage, 11:59 am, Jan 7, set the thermal properties cycle to 10 minutes (from 180 minutes) so that Steve/Cody could check the voltages. At 12:05, did the same at lake. Steve determined at sage that the variation seen on Vheat is due to the charging system. At 1:30 pm, changed sage back to 180 minute cycle. At 1:32, changed lake back. At 1:37pm, changed grass to 10 minutes. At 3:51pm, changed grass back to 180 minutes. The TP instructions are numbers 68 and 72 The values for normal 3 hour cycling are: MODE 01:0000 68\r\n 68:P92 \r\n 01:0000 \r\n 02:0180 \r\n 03:30 \r\n \r\n MODE 01:0000 72\r\n 72:P92 \r\n 01:0003 \r\n 02:0180 10\r\n 03:30 \r\n \r\n For 10 minute cycling, change the second parameter from 180 to 10.
The logger at lake was looping through the thermal properties code every 3 minutes, instead of every 3 hours. The code was changed to 3 hour looping at 11:30am, Dec 19
The Sage, Lake and Grass C soil samples were not treated the same as were 
the A and B. Rather, the cores were extracted in toto and will be used for 
bulk density determination. Because of the limitation of the balance the cores
were broken in two. The core portions were put into larger Al container, were
weighed and then left to dry until the next visit.
In all cases the ~20 cm of snow was removed and the ice crust atop the soil 
surface was broken and removed. At Grass dead grass stalks were also removed.
Sage C            Tare               Fresh(18 Dec)           Dry(1/7/03)
0 - 8cm  ZI        8.0               143.2		     122.0
         ZII       8.0               149.0		     115.2
Lake C            Tare               Fresh(18 Dec)           Dry(1/7/03)
0 - 8 cm ZIII      8.0               120.5		     87.0
         ZIV       8.0               168.8		     127.2
Grass C           Tare               Fresh(18 Dec)           Dry(1/7/03)
0 - 8 cm ZV        8.0                76.2  top 4 cm peaty   34.3
         ZVI       8.0               136.1  bottom 4 cm      106.4
Weighed the soil samples collected on Mon 16 Dec (see log entry 67)
11:40 Sage site. .At B, dug away 10 cm of snow. There was a 2 -3 cm ice crust 
on the soil surface. Very hard going to sink the core as the ground was frozen.
Although the snow was completely removed from the surface before coring the ice 
incorporated in the soil melted in the base to yield mud.
The corer stem broke and had to be replaced with a pice of pipe.
                   Fresh        Dry (18 Dec)
   0-3 cm  AI      25.3         17.6
           AII     20.9         14.7
           AIII    21.9         16.1
   3-8 cm  BI      26.1         21.6
           BII     22.6         18.8
           BIII    22.3         18.8
13:30 Lake site At B dug down through 30 cm of snow and brushed soil surface
clear of snow
                   Fresh        Dry (18 Dec)
  0-3 cm  CI       13.3         11.5
          CII      16.6         14.4
          CIII     14.6         12.5
  3-8 cm  DI       13.9         11.3
          DII      15.7         12.7
          DIII     19.1         15.3
14:00 Grass site. At B 25 - 30 cm of snow. Dug down and then broke up crust
incorporating dead grass stems and ice. Cleaned this away and started core
at peaty surface.
                   Fresh       Dry (18 Dec)
  0-3 cm  EI        8.2         2.5
          EII       7.4         2.4
          EIII      6.7         2.4
  3-8 cm  FI       15.9        12.3
          FII      17.3        12.7
          FIII     19.7        15.3
Gravimetric soil samples are to be:
                                   top  3 cm
                                   next 5 cm
The plan is to do Sage A, Lake A and Grass A on Mon, 16 Dec, dry overnight
 ......................B.......B           B    Tue  17..................
.......................C ......C.......... C....Wed..18    leave in oven
Also to take total cores from three sites for bulk density and leave in oven
11:20 Sage site. Approx 10 cm of snow between bushes with almost no snow under some bushes.At A, under a bush with no snow cover, core sample.
                   Fresh        Dry (17 Dec)
   0-3 cm  AI      16.6         11.7
           AII     12.4          8.5
           AIII    13.1          8.8
   3-8 cm  BI      22.7         20.6
           BII     23.7         21.6
           BIII    24.2         21.9
12:00 Lake site Fairly uniform 20 -25 cm snow. Couldnt locate A so went approx 
the correct distance from box. Dug down through snow to soil and brushed the surface clear.
                   Fresh        Dry (17 Dec)
  0-3 cm  CI       15.6         11.0
          CII      14.5         10.4
          CIII     17.1         12.5
  3-8 cm  DI       23.4         21.5
          DII      25.8         23.8
          DIII     25.4         23.2
12:40 Grass site. Grass sticking out of 0-8 cm snow.
Note that the top 1 cm was grass stalks. I discarded this and separated the dirtr component from the next 3 cm. The lower 5 cm was mostly dirt.
                   Fresh       Dry (17 Dec)
  0-3 cm  EI       5.5          2.9
          EII      4.6          2.4
          EIII     5.0          2.7
  3-8 cm  FI       9.5          6.8
          FII     11.3          8.3
          FIII    12.7          9.2
TP01 SN 200235 installed at lake site ~1645. It is installed about at 5cm in the frozen sand. It still has an air gap, though somewhat neater (and thus smaller) than at the tower site -- I'm learning.
TP01 SN 200234 installed at tower site ~1345. (It was connected an hour earlier, but it was on top of the snow.) It is installed about at 5cm in the (peat-like) dirt. I'm sure that there is an air-gap beside it, but hopefully this will heal with time.
Dec 12, 16:45 MST As in logbook entry #51 at the grass site, changed the CR10x logger code at lake and sage so that there is a 10 msec delay before sampling the Tsoil probes.
Dec 11, 10:27 MST: Added 10 millisecond delay to P11 instruction (#50) which is used to sample T107 Tsoil probes. This is the 03 parameter, which we changed from 1 to 21. Lab testing showed that this reduces noise, and so we hope this will reduce the noise on Tsoil that we see when the radiometer fans are on. This changed reduced the noise significantly, except on the 10cm probes, which are sampled at another point in the CR10 code. Dec 11, 15:13 MST Changed the P11 instruction for the 10cm probes too (instruction 60). Here is the rserial output: MODE 01:0000 60\r\n 60:P11 \r\n 01:01 \r\n 02:05 \r\n 03:01 \r\n 04:0027--\r\n 05:+1.0000 \r\n 06:+0.0000 \r\n MODE 01:0000 60\r\n 60:P11 \r\n 01:01 \r\n 02:05 \r\n 03:01 21\r\n 04:0027--\r\n 05:+1.0000 \r\n MODE 01:0000 60\r\n 60:P11 \r\n 01:01 \r\n 02:05 \r\n 03:21 \r\n 04:0027--\r\n 05:+1.0000 \r\n 06:+0.0000 \0x7f\r\n Oops: looks like I didn't restart the logger with a "*0" and it didn't resume until I rserial'd to it again at about 16:20.
the soil profile array interface cable was installed on Wed, Dec 4th we are now receiving data
On Thurs, 10 Oct''02 Tony and John visited the FLOSS site
The Echo soil sensors at Sage, Lake and Grass were reinstalled.
They had been installed horizontally the previous week and had been watered 
to help them meld into the soil. This melding was apparently sucessful at 
Sage and Lake, where the soil was suitable, but not so sucessful at Grass,
where the soil was actually peat.
The Echo probes were reinstalled vertically, with their midlines at the 
designated depths of 5 cm and 10 cm. They were positions at right angles
to each other with ~12cm between their cabled ends.
Sage  A 1 @  5cm
      A 2 @ 10cm
Sage  B 3 @  5cm
      B 4 @ 10cm
Sage  C 5 @  5cm
      C 6 @ 10cm
Each set of Echo probes waterd with 1 gallon
Lake  A 1 @  5cm
      A 2 @ 10cm
Lake  B 3 @  5cm
      B 4 @ 10cm
Lake  C 5 @  5cm
      C 6 @ 10cm
Each set of Echo probes watered with 1 gallon
 
Grass A 1 @  5cm (It is possible that A1 is actually at 10cm)
      A 2 @ 10cm (It is possible that A2 is actually at  5cm)
Grass B 3 @  5cm
      B 4 @ 10cm      
Grass C 5 @  5cm
      C 6 @ 10cm
Each set of Echo probes watered with 2 gallons
     
10/10/02 The sensors apparently didn't have the white dot and were installed up-side down. This is very apparent in the data. Fix: can either be swap wires in logger, or change sign of logger coef.
10/10/02 Soil Echo Probes Rotated: from Horizontal to Vertical orientation at Sage/Lake/Grass. Water was dumped on to help bed the sensors in. Tony said they already were looking good in that respect. The logger cable was added to the Lake site. Both Sage/Lake are now reporting data through goes.
On 2 Oct''02 the locations for the soil sensors for Grass, Sage and Lake soil sensor arrays were chosen by Tony, Steve O and Jielun. Tony installed the sensors at Grass and Sge on 2 Oct and installed the sensors at Lake on the 3 Oct. After installation the sensors were wetted down with approximately 6 liters of water to "bed them in". It is anticipated that given another month of warm weather the local soil moisture will come to equilibrium before the start of the program in December. For the schematic of the placement see